I need help in finding the equation that relates C AND X FROM THIS PROBLEM 100 ITEMS WILL COST $10,000 TO PRODUCE AND 300 ITEMS WILL COST $22,000 TO PRODUCE
Answer by tutorcecilia(2144)

Since the cost depends upon the number of items, let the cost equal the y value and number of items equal the x value:
x=100
y=$10,000
(100; 10000)
.
x=300
y=$22,000
Point (300; 22000)
.
Find the slope of the line using the two points.
m=(22,000-10,000)/(300-100)=12,000/200=60
Find the equation of the line using one of the formulas for a line.
y=mx+b
22,000=(60)(300)+b [Plug-in the values and solve for b]
22,000=18000+b
22,000-18,000=b
4,000=b [Plug-in the slope and y-intercept value]
y=60x+4,000
C(x)=60x+4,000